hey iv a freashman and my teacher is makeing us do a sheet a sertain way can u show me all the ways to solve (x-8=0) (6x+5=13) (67=x-65) (x-4=9) (6=x+2) (5b=145) (-4=(2/5a) (7y=28) (m/7=12) im kinda in need of help cuz she wants me to do it in a certain way and if i dont ill get all of them wrong even though i got them right so plz help
Answer by checkley77(12844)

(x-8=0) ADD 8 TO BOTH SIDES.
X-8+8=0+8
X=8 ANS.
--------------
(6x+5=13) SUBTRACT 5 FROM BOTH SIDES.
6X+5-5=13-5
6X=8 DIVIDE BOTH SIDES BY 6.
6X/6=8/6
X=8/6 REDUCE THE 8/6
X=4/3 ANS.
---------------------------------
(67=x-65) ADD 65 TO BOTH SIDES
67+65=X-65+65
132=X ANS.
----------------------------------
(x-4=9) ADD 4 TO BOTH SIDES.
X-4+4=9+4
X=13 ANS.
--------------------------------
(6=x+2) SUBTRACT 2 FROM BOTH SIDES.
6-2=X+2-2
4=X ANS.
---------------------------
(5b=145) DIVIDE BOTH SIDES BY 5.
b=145/5
b=29 ans.
--------------------------------
(-4=(2/5a) cross multiply
-4*5a=2
-20a=2 divide both sides by -20
-20a/-20=2/-20
a=-1/10 ans.
----------------------------------
(7y=28) divide both sides by 7.
7y/7=28/7
y=4 ans.
---------------------------------------
(m/7=12)cross multiply.
m=7*12
m=84 ans.